OpenStudy (anonymous):

a certain compound of oxygen and sulfur has a molar mass of 80. what is the empircial formula?

OpenStudy (anonymous):

I think it's either \[SO_3\] or \[S_2O\] I'm not really sure how to solve for which one it could be though.

OpenStudy (anonymous):

The way I did it was to find a ratio between the compounds molar mass and individual molar masses of the elements: \[\frac{ 80 }{ 16 }=5, \frac{ 80 }{ 32.065 }=2.5\] So you can have a maximum of 2 sulfurs, then you subtract the molar mass of sulfur and divide by the molar mass of oxygen, which is why I can get two answers. \[80-32=48 then \frac{ 48 }{ 16 }=3 Oxygens\] \[80-(2*32)=16 then \frac{ 16 }{ 16 }=1 Oxygen\]